### v1

# Towards Next-Generation Recommender Systems: A Benchmark for Personalized Recommendation Assistant with LLMs

Abstract.
Recommender systems (RecSys) are widely used across various modern digital platforms and have garnered significant attention. Traditional recommender systems usually focus only on fixed and simple recommendation scenarios, making it difficult to generalize to new and unseen recommendation tasks in an interactive paradigm. Recently, the advancement of large language models (LLMs) has revolutionized the foundational architecture of RecSys, driving their evolution into more intelligent and interactive personalized recommendation assistants. However, most existing studies rely on fixed task-specific prompt templates to generate recommendations and evaluate the performance of personalized assistants, which limits the comprehensive assessments of their capabilities. This is because commonly used datasets lack high-quality textual user queries that reflect real-world recommendation scenarios, making them unsuitable for evaluating LLM-based personalized recommendation assistants. To address this gap, we introduce RecBench+, a new dataset benchmark designed to access LLMs’ ability to handle intricate user recommendation needs in the era of LLMs. RecBench+ encompasses a diverse set of queries that span both hard conditions and soft preferences, with varying difficulty levels. We evaluated commonly used LLMs on RecBench+ and uncovered below findings: 1) LLMs demonstrate preliminary abilities to act as recommendation assistants, 2) LLMs are better at handling queries with explicitly stated conditions, while facing challenges with queries that require reasoning or contain misleading information. Our dataset has been released at https://github.com/jiani-huang/RecBench.git.
Recommender System, Recommendation Assistant, Large Language Models (LLMs), Dataset Benchmark.
†Correspondence to: Wenqi Fan and Qing Li, Department of Computing, The Hong Kong Polytechnic University.
copyright: acmlicensedjournalyear: 2018isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Illustration of Different Recommendation Paradigms. (a) Traidtional-based recommendations generate suggestions in simple and fixed scenarios by learning user/item representations, like Top-K or similar item recommendations. (b) LLM-based recommendation relies on fixed and simple query templates, restricting users to predefined formats. (c) In contrast, our benchmark evaluates the LLM’s ability to handle complex, flexible user queries as a recommendation assistant.
Recommender Systems (RecSys) play a crucial role in modern digital platforms, enabling personalized content delivery across domains such as e-commerce (Chen et al., 2023), entertainment (Gao et al., 2023), and education (Urdaneta-Ponte et al., 2021). For example, YouTube once reported that 60% of the clicks on its homepage came from recommended content (Davidson et al., 2010), revealing the huge commercial value of RecSys. Generally, these systems are trained to handle typical recommendation scenarios, such as “Customers Who Viewed This Also Viewed” to recommend alternatives or similar items and “Based on Your Browsing History” to recommend Top-K items of interest using a user’s historical data, as shown in Figure 1 (a). As one of the most representative algorithms, matrix factorization (MF) techniques (Koren et al., 2009) are proposed to learn latent representations of users and items based on user-item interactions for predicting matching scores in recommendations. However, these traditional recommendation methods are applicable only to a limited range of simple and fixed recommendation scenarios, relying on pre-designed and well-trained recommendation methods, and they struggle to generalize to new and unseen recommendation tasks (i.e., recommendation queries) in an interactive paradigm. For instance, if a user searches for ‘a durable laptop for graphic design under \$1500’, traditional approaches like MF depend only on the correlation between users and items, failing to capture these nuanced requirements and deliver appropriate recommendations.
Recently, large language models (LLMs) like GPT-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and LLaMA (Dubey et al., 2024) have demonstrated strong capabilities in generalization and reasoning. For instance, AUTOGLM (Liu et al., 2024b) acts as an autonomous assistant, leveraging the LLMs’ human-like intelligence in understanding and planning capabilities to help users control digital devices via GUIs and address complex tasks such as booking flights (Singh et al., 2024), collecting data (Schmidgall et al., 2025), and coding (Kazemitabaar et al., 2024; Pinto et al., 2024), significantly enhancing convenience in people’s daily lives. Inspired by the powerful capabilities of LLMs, many efforts have been made to incorporate LLMs into RecSys as personalized recommendation assistants in an interactive paradigm. For example, LLaRA (Liao et al., 2024) directly transforms users’ historical interactions into natural language formats (i.e., prompts) for personalized recommendations. Benefiting from LLMs, as shown in Figure 1 (b), these personalized recommendation assistants can handle more interactive and challenging scenarios, such as “Tell You What You Like,” which allows users to ask questions in a conversational manner and receive corresponding recommendations.
Despite the remarkable success, current studies lack a comprehensive evaluation of the recommendation capability of these LLM-based personalized recommendation assistants. Specifically, in the phases of both training and testing, most existing methods leverage fixed and simple task-specific prompt templates to generate recommendation responses, and assess the recommendation performance of these assistants, such as “Will the user like {movie_i}. Please answer Yes or No.” to determine whether to recommend a specific item, or “Please choose a movie from the candidate list {movie_1, movie_2, … }” to recommend a user-liked item. However, as shown in Figure 1 (c), this testing paradigm for evaluation fails to align with real-world practical scenarios, where the user queries can be highly complex when they interact with the LLM-based recommendation assistant, such as “Recommend me a funny romantic movie like Alien, starring Leonardo Wilhelm DiCaprio, and suitable for female audiences.” In this case, the recommendation assistant must not only understand the user’s desire to find movies starring Leonardo Wilhelm DiCaprio, but also filter out funny romantic movies that are appropriate for female viewers. Moreover, the recommendation assistant should be robust to perturbations in user input, i.e., it should recognize that the user is looking for a funny animated movie rather than a horror film like Alien. This requires the assistant to deeply comprehend multiple constraints within the user’s input and provide suitable and robust recommendations through remarkable reasoning capability, which is quite challenging. However, promising results on the existing benchmark datasets may not be extended to the new recommendation paradigm in the era of LLMs.
Existing frequently used datasets, such as LastFM (Schedl, 2016), Amazon Beauty (Ni et al., 2019), and Movielens-1M (Harper and Konstan, 2015), typically only include basic user information, item metadata (e.g., title, categories, and descriptions), and user-item interactions. These datasets are more suitable for assessing the accuracy performance of traditional RecSys but are not appropriate for effectively evaluating recommendation assistants in the era of LLMs, due to the lack of high-quality textual user queries which can meet practical recommendation scenarios.
To bridge this gap, we introduce a new benchmark, RecBench+, tailored for evaluating the potential of LLMs to act as personalized recommendation assistants. Specifically, we propose a new dataset comprising approximately 30,000 high-quality and complex user queries, constructed based on multiple frequently-used recommendation datasets (e.g., Movielens-1M (Harper and Konstan, 2015) and Amazon book [^1] https://archive.ics.uci.edu/ml/datasets/Amazon+book+reviews). These queries vary in difficulty, number of conditions that user specified, and user profiles, effectively reflecting the diverse user needs in real-world scenarios. We comprehensively evaluate the performance of seven widely used LLMs on RecBench+, including GPT-4o (Hurst et al., 2024), Gemini-1.5-Pro (Team et al., 2024), DeepSeek-R1 (Guo et al., 2025), etc. We further analyze various factors that impact model performance, such as the number of conditions and the integration of user interaction history. Through extensive experiments with state-of-the-art LLMs in our benchmark, we obtain several insightful findings: (1) LLMs demonstrate preliminary abilities to act as recommendation assistant, but their strengths vary. GPT-4o and DeepSeek-R1 excel at user queries with clear conditions, while Gemini-1.5-Pro and DeepSeek-R1 perform better in queries without clear conditions but require a clear understanding of user profiles. (2) For queries with user-specified conditions, LLMs are more adept at handling those with explicitly stated conditions. In contrast, queries that require reasoning or contain perturbations pose greater challenges. However, models equipped with advanced reasoning capabilities tend to perform better on such tasks. (3) For queries requiring user profile understanding, LLMs show varying performance across different user interests and demographics. For example, LLMs may be better at providing recommendations for popular interests and tend to perform more effectively for female users than for male users. Our contributions are summarized as follows:
  • Novel Recommendation Paradigm: We introduce a new paradigm for next-generation recommender systems, exploring the effectiveness of LLMs acting as personalized recommendation assistants in an interactive and intelligent way. As the first benchmark to effectively integrate LLMs in this interactive and intelligent manner, we transform the traditional recommendations paradigm into a more personalized and context-aware experience, setting a new standard for the field.
  • Dataset Construction: We present a comprehensive dataset called RecBench+ to simulate practical recommendation scenarios in the era of LLMs. Our dataset encompasses approximately 30,000 high-quality and complex user queries that vary in difficulty, the number of conditions specified by users, and user profiles, reflecting the diverse user needs in real-world recommendation scenarios. To the best of our knowledge, this is the first public dataset that can be used to effectively evaluate the performance of personalized recommendation assistants in the era of LLMs.
  • Comprehensive Evaluation: We conduct extensive experiments with state-of-the-art LLMs, analyzing their strengths and limitations in this task and providing actionable insights.
2. The RecBench+ Benchmark
In this section, we first present an overview of the RecBench+ benchmark. Then, we provide the construction methods of each data type included in our proposed dataset.
2.1. Overview
In practical applications, user demands can generally be categorized into two types, each requiring specialized capabilities from LLM-based recommendation assistants. The first type is driven by specific conditions or constraints, such as asking for movies by a particular director or within a certain genre. This scenario demands strong reasoning capabilities from the recommendation assistant, as it requires the assistant to accurately interpret and apply the given constraints to generate appropriate recommendations. The second type does not involve specified conditions but requires recommendation assistants to generate recommendations based on an understanding of the user’s profile (e.g., user interest, occupation, gender), which emphasizes the assistant’s ability to interpret the profile and provide personalized suggestions. Given that these two scenarios demand different capabilities from the recommendation assistant, we construct two distinct query types to evaluate their performance: Condition-based and User Profile-based. Below, we introduce the two query categories in detail.
2.1.1. Condition-based Query
Condition-based Query simulates scenarios where users have a clear understanding of what they want and set concrete conditions for the expected items. For example, in a movie recommendation assistant, a user might seek films featuring specific directors, actors, genres, or other attributes. The core challenge here lies in correctly understanding and satisfying all constraints in the query, particularly when queries include multiple conditions. To comprehensively evaluate the recommendation assistant’s reasoning ability to fully interpret and fulfill constraints, we categorize Condition-based Query into three types according to how users express their needs when seeking recommendations. Typically, users explicitly state their conditions for the desired item, which we define as Explicit Condition Query. In addition to clearly stating the conditions, users may only provide vague or incomplete information about the items they need, requiring the recommendation assistant to infer the underlying conditions from the context, which we regard as Implicit Condition Query. Additionally, users might have mistaken information or make typos for the item, leading to Misinformed Condition Query. The following are detailed explanations:
  • Explicit Condition Query:
    These queries explicitly specify the conditions that the recommended items should satisfy in the natural language format. For example, the query, “Can you recommend movies featuring Gwyneth Paltrow?”, directly states the condition (“featured by Gwyneth Paltrow”) the recommendations must satisfy. Explicit condition queries are relatively straightforward because the constraints are clearly articulated.
  • Implicit Condition Query: Unlike Explicit Condition Query, Implicit Condition Query require the recommendation assistant to infer the conditions of the target items, because the user may not provide complete or specific details about the items. For example, “Can you suggest movies with the same cinematographer as Stay Hungry (1976) and Beethoven (1992)?” does not mention the cinematographer’s name (i.e., David Worth), forcing the recommendation assistant to conduct in-depth reasoning about the name of the cinematographer first. Implicit condition queries aim to assess assistants’ reasoning and knowledge retrieval capabilities, making them more challenging than explicit condition queries.
  • Misinformed Condition Query: Users may sometimes provide incorrect or misleading information about items in their queries. For example, the query, “I like ‘Avatar’ directed by Spielberg. Please recommend similar movies directed by him.”, incorrectly attributes the movie Avatar to Steven Spielberg. It is expected that recommendation assistants can first identify and correct this misinformation–recognizing that Avatar was not directed by Spielberg—before generating relevant recommendations. This subcategory presents a significant challenge to the robustness of the recommendation assistant, requiring it to accurately identify and filter out inaccurate information in users’ queries.
2.1.2. User Profile-based Query
In many real-world scenarios, users may not provide specific conditions to filter items but instead expect to be recommended personalized results based on their profile information, such as interests (e.g., interest in special effects) or demographic attributes (e.g., gender, occupation). This is because users often lack a clear understanding of their own needs and instead seek recommendations that intuitively align with their background or preferences. To address this common situation, we introduce User Profile-based Query, which simulates these scenarios by requiring the recommendation assistant to generate personalized suggestions based on user profile information. We further divide User Profile-based Query into Interest-Based and Demographics-Based Query, since interests reflect users’ dynamic preferences, while demographics capture more stable attributes like gender or occupation that influence their needs.
  • Interest-based Query: These queries reflect user interests that can’t be clearly defined by specific conditions. For example, users who watched “The Tree of Life” and “2046” in history may prefer movies with unique storytelling and visual beauty. To capture such interests, we create queries like, “I just watched ‘The Tree of Life’ and loved its poetic storytelling and visual design. Can you recommend other movies with a similar tone?” Such queries don’t have explicit conditions but require the personalized recommendation assistant to understand and suggest items based on the user’s personal interest.
  • Demographics-based Query: These queries focus on how users’ demographic attributes influence recommendations. Different users, based on their age, gender, or occupation, may have distinct preferences and needs. For example, philosophy professors might frequently be interested in films that explore moral dilemmas. By leveraging these patterns, we can craft queries such as: “I’m a philosophy professor looking for movies with complex themes and moral dilemmas. Can you recommend any?” By aligning with real-world demographic behaviors, these queries can mirror the personalized recommendations that a personalized recommendation assistant should provide, ensuring that the suggestions are relevant and tailored to the user’s background and identity.
2.1.3. Statistics of the Constructed Benchmark Dataset RecBench+
The benchmark dataset’s statistical breakdown is shown in Table 1, illustrating the distribution of query types. The movie dataset is sourced from Movielens-1M [^2] MovieLens-1M provides detailed user demographic information that can be used for demographics-based queries, while other versions, such as MovieLens-10M and MovieLens-20M, lack such information. Therefore, we only can adopt the MovieLens-1M dataset to construct our dataset., while the book dataset is from Amazon-Book. We then describe the construction details of Condition-based Query and User Profile-based Query.
Table 1. Statistics of RecBench+
Major Category Sub Category Condition Movie Book
Condition-based Query Explicit Condition 1 2,225 2,260
2 2,346 2,604
3,4 426 271
Implicit Condition 1 1,753 1,626
2 1,552 2,071
3,4 344 213
Misinformed Condition 1 1,353 1,626
2 1,544 2,075
3,4 342 215
User Profile-based Query Interest-based - 7,365 2,004
Demographics-based - 279 0
Queries in Total 19,529 14,965
Number of Users 6,036 4,421
Number of Items 3,247 9,016
Number of User-Item Interactions 33,291 29,285
2.2. Condition-based Query Construction
Figure 2. An overview of the pipeline of constructing Condition-based Query. The process starts by constructing an Item Knowledge Graph (KG) that links items to their attributes. The second step is to identify item sets from the user’s history that share the same relations (e.g., movies directed by the same person) using the KG. These shared relations serve as the basis for constructing different types of conditions for the query. Finally, we use LLMs to simulate users and generate queries based on conditions, and the aforementioned item set serves as the ground truth.
User queries often express preferences through conditions corresponding to item attributes, such as filtering movies by genre, director, or actor. However, due to the intricate relationships between items and their attributes, it is challenging to generate realistic, diverse queries that reflect how users combine these attributes. It is worth noting that Knowledge Graph (KG), as a structured representation, explicitly connects items (e.g., movies) to their attributes (e.g., directors) and captures interdependencies between them (e.g., “director X frequently collaborates with actor Y”). Therefore, we propose leveraging KG to extract meaningful conditions and generate queries aligned with real-world user behavior. Specifically, our approach contains three key steps: Item Knowledge Graph (Item KG) construction, shared relation extraction, and query generation, as shown in Figure 2. We first build the Item KG, which links items to their attributes. Next, we identify item sets from the user’s interaction history that share the same attributes (e.g., movies directed by Christopher Nolan), forming the conditions for the query. Finally, we adopt LLMs to simulate users and generate queries based on these extracted conditions. The item sets that satisfy the conditions can serve as the ground truth for evaluation.
2.2.1. Item KG Construction
The Item KG forms the basis for generating realistic Condition-based Query in RecBench+. We take advantage of different KGs for the movie and book domains:
  • Movies: We extract data from Wikipedia, focusing on 7 key attributes like directors, actors, composers, and genres. Each movie node is linked to these attributes.
  • Books: We use metadata from the Amazon Book Dataset, connecting attributes like authors and categories to book nodes.
Once the KGs are built, we combine them with traditional recommendation datasets, Movielens-1M and Amazon-Book, to generate conditions and queries step by step.
2.2.2. Shared Relation Extraction
Shared relations are common attributes found among items in a user’s interaction history. For example, if a user has watched multiple movies by the same director or starring the same actor, these shared attributes help generate Condition-based Queries. Given a user $u$ with an interaction history $\mathcal{H}_{u}=\{i_{1},i_{2},\ldots,i_{k}\}$, we employ a KG retrieval function $\mathcal{R}$ to identify shared attributes (relations) across subsets of items in $\mathcal{H}_{u}$. Each shared relation is defined as a tuple $(r,t)$, where $r$ is the type of relation (e.g., “directed by”, “starring”) and $t$ is the target value (e.g., name of director). The extraction process produces multiple groups of shared relations and their corresponding subsets of items, which can be represented as follows:
$\mathcal{R}(\mathcal{H}_{u},KG)=\left\{\left(\mathcal{G}_{\text{sub}},\mathcal{C}_{\text{shared}}\right)\;\middle|\;\mathcal{G}_{\text{sub}}\subseteq\mathcal{H}_{u},\right.$
(1) $\left.\mathcal{C}_{\text{shared}}=\{(r_{1},t_{1}),(r_{2},t_{2}),\ldots\}\right\}.$
Here, $\mathcal{G}_{\text{sub}}$ represents a subset of items sharing the same relations $\mathcal{C}_{\text{shared}}$. The extracted shared relations will later serve as conditions for query generation.
2.2.3. Query Generation
After extracting shared relations from the user’s interaction history, we generate three types of Condition-based Query: explicit, implicit, and misinformed. The steps to generate these conditions and final queries are as follows:
  • Explicit Condition Construction: For explicit conditions, we directly adopt the shared relations extracted from the user’s interaction history. Specifically, we set the explicit condition $\mathcal{C}_{\text{explicit}}\ =\mathcal{C}_{\text{shared}}$. This means the conditions are directly derived from the shared attributes (e.g., director, genre) associated with the items the user has interacted with.
  • Implicit Condition Construction: For implicit conditions, instead of directly stating the specific value (e.g., the director’s name), we leverage the KG to describe them indirectly through related items. For example, instead of directly referencing James Cameron, we might use “the director of ‘The Abyss’ ”. Specifically, for a shared relation $(r_{m},t_{m})\in\mathcal{C}_{\text{shared}}$ (e.g., (director, Cameron)), we replace the value that needs to be inferred, $t_{m}$ (e.g., Cameron), with an indirect reference that describes $t_{m}$’s relation with another item $i_{k}$ (e.g., The Abyss) from the KG. Mathematically, this can be expressed as:
    (2) $i_{k}\in\{i\in\mathcal{I}\mid(t_{m})\xrightarrow{r_{m}}(i)\text{in KG}\},$
    where $\mathcal{I}$ is the set of all items in the KG. The resulting implicit condition is then:
    (3) $\mathcal{C}_{\text{implicit}}=(r_{i},\text{ref}(i_{k})),$
    where $\text{ref}(i_{k})$ denotes a textual reference to $i_{k}$. For example, a shared relation like (director, Cameron) can be transformed into an implicit condition (director, “the director of ‘The Abyss’ ”), where “The Abyss” ($i_{k}$) is retrieved from the KG as being connected to “Cameron” ($t_{0}$) via the “director” relation ($r_{0}$).
  • Misinformed Condition Construction: For misinformed conditions, we intentionally introduce factual error into the conditions, simulating real-world user errors where they might mistakenly reference attributes or relationships. To generate a misinformed condition, for a shared relation $(r_{i},t_{i})\in\mathcal{C}_{\text{shared}}$, we randomly select one or more items $\{i_{1},i_{2},\ldots,i_{m}\}\in\mathcal{I}$ from the KG that do not have the specified relationship $r_{i}$ with $t_{i}$ (i.e., $i_{k}\not\longrightarrow t_{i}$ for each $i_{k}\in\{i_{1},i_{2},\ldots,i_{m}\}$). We then construct the condition with “error info” that falsely claims these items are related to $t_{i}$ through $r_{i}$:
    (4) $\mathcal{C}_{\text{misinformed}}=(r_{i},t_{i},\text{error info: }\{i_{1},i_{2},\ldots,i_{m}\}\xrightarrow{r_{i}}t_{i}),$
    where the set $\{i_{1},i_{2},\ldots,i_{m}\}$ contains items from the KG that do not actually share the relationship $r_{i}$ with $t_{i}$. For example, given a condition (director, Cameron), a misinformed condition might be (director, Cameron, error info: Cameron is the director of ‘Startreck’).
  • Query Generation: For each condition type, we apply GPT-4o (Hurst et al., 2024) to generate corresponding queries:
    (5) $q=LLM(\mathcal{C},\mathcal{P}),$
    where $\mathcal{P}$ is the prompt, see appendix H. For each generated query $q$, we construct a sample consisting of:
    (6) $(q,\mathcal{H}_{u}^{\prime},\mathcal{G}_{q}),\ \text{where:}\ \mathcal{H}_{u}^{\prime}=\mathcal{H}_{u}\setminus\mathcal{G}_{q},$
    (7) $\quad\mathcal{G}_{q}=\{i\in\text{KG}\mid\mathcal{C}_{\text{shared}}\subseteq\mathcal{A}_{i}\}.$
    Here, $\mathcal{H}_{u}^{\prime}$ is the user’s interaction history excluding items in $\mathcal{G}_{q}$, and $\mathcal{G}_{q}$ contains the ground truth items satisfying the query conditions. $\mathcal{A}_{i}$ represents the attributes of item $i$ in the KG.
2.3. User Profile-based Query Construction
Figure 3. An overview of the pipeline of constructing User Profile-based Query. (a) For Interest-based Query (top), we identify common interests by extracting collaborative item sets and preceding item sequences from user interactions, and queries are generated based on inferred interests. (b) For Demographics-based Query (bottom), users are grouped by demographics, and queries are generated using demographics and popular items within each group.
In addition to providing specific constraints, users may provide only basic personal information such as watching history or demographic attributes (e.g., age, occupation). To evaluate the LLM recommendation assistant’s ability to handle such scenarios, we introduce User Profile-based Query, which aims to generate personalized recommendations based on user profile information without requiring specific filtering conditions. As detailed in Section 2.1.2, User Profile-based Query contains two subcategories: Interest-based Query and Demographics-based Query. The construction process is illustrated in Figure 3 and we describe the construction process below.
2.3.1. Interest-based Query
These queries are designed to capture user interests that are not explicitly defined by specific conditions but are inferred from user interactions. To effectively simulate real-world scenarios, we focus on capturing the collective interests that emerge from shared behaviors from multiple users, which are more representative of general preferences and trends. For example, if many users who watched “The Dark Knight” also watched “Inception”, this suggests a clear interest in action and suspense. In this way, we can craft queries that reflect collective user interests derived from shared behaviors. Specifically, we first identify Common Interest Set, which are sets of items that frequently interact consecutively by multiple users. Let $\mathcal{H}_{u}$ denote the interaction history of user $u$, and =$\mathcal{H}=\{\mathcal{H}_{u}\mid u\in\mathcal{U}\}$ denote the set of all users’ interaction histories. The common interest set can be defined as:
(8) $S=\{s\mid\exists u\in\mathcal{U},s\subseteq\mathcal{H}_{u},f(s)\geq\theta\},$
where $f(s)$ is the frequency of sequence $s$ across all users’ histories, and $\theta$ is a predefined frequency threshold.
Once common interest sets are identified, we proceed to extract the preceding item sequences—that is, the sequences of items that commonly appear before the items of the common interest set. Let $P(s)$ denote the set of preceding item sequences for target sequence $s$. The preceding sequences can be defined as:
(9) $P(s)=\{p\mid\exists u\in\mathcal{U},p\prec s\subseteq\mathcal{H}_{u}\},$
where $p\prec s$ denotes that sequence $p$ appears immediately before sequence $s$ in the interaction history $\mathcal{H}_{u}$.
Based on the identified common interest set and preceding item sequences, we employ an LLM to analyze the patterns and infer the reasoning behind multiple users’ common interests. The generated query aims to reflect both the contextual features of the items and the common interests of users. The prompt used to generate these queries is provided in Appendix H.
2.3.2. Demographics-based Query
Individuals with similar demographic backgrounds, like age, gender, or occupation, often exhibit similar preferences and interaction patterns. For example, kids typically enjoy animated movies like Disney films, while teenagers may prefer action and superhero movies. Therefore, by grouping users who share the same demographic attributes (e.g., students aged 18–24), we can effectively construct recommendation queries influenced by demographic attributes. Specifically, we categorize users into different user groups based on permutations and combinations of demographic attributes. For each user group, we first identify the set of items most frequently consumed by users within that group. Subsequently, we provide both the user group demographics and the list of most popular items as input to an LLM. The LLM then analyzes the underlying patterns of each user group. Subsequently, LLM generates a query that encapsulates both the distinct characteristics of the user group demographics and the preferences reflected in the most popular items. This query is designed to test the ability of recommendation assistants to provide personalized suggestions based on demographic attributes. The list of most popular items will serve as the ground truth for evaluating the relevance of recommendations generated in response to the query.
3. Experiments
3.1. Experiments Settings
3.1.1. Baselines
We evaluate several widely used models, including GPT-4o (2024-08-06) (Hurst et al., 2024), GPT-4o-mini (2024-07-18) (Hurst et al., 2024), Llama (Llama-3.1-70B-Instruct) (Dubey et al., 2024), Gemini (gemini-1.5-pro-002) (Team et al., 2023), Claude (claude-3-5-sonnet-20241022), DeepSeek-V3 (Liu et al., 2024a), and DeepSeek-R1 (Guo et al., 2025). However, we were unable to test the GPT-o1 (Jaech et al., 2024) model from OpenAI because it flagged our prompts as violating its usage policy.
3.1.2. Evaluation Metrics
We utilize the below metrics to evaluate LLM performance. To simulate real-world scenarios where users typically do not predefine the number of recommendations they need, we do not specify a fixed number of recommendations K in the testing prompts; however, we also include several experiments with a fixed K in the Appendix F for reference.
Table 2. Performance on Condition-based Query. The  /  means the best/the second-best result. In Misinformed Condition Query, a higher FTR rate is better as it shows the model can reject recommendations when queries have factual errors.
Domain Model Explicit Condition (Easy) Implicit Condition (Medium) Misinformed Condition (Hard) Average
[…]
Movie GPT-4o-mini 0.185 0.322 0.531 0.009 0.083 0.167 0.198 0.017 0.028 0.060 0.153 0.104 0.099 0.183 0.294
GPT-4o 0.308 0.408 0.714 0.016 0.145 0.224 0.301 0.021 0.019 0.039 0.106 0.270 0.157 0.224 0.374
Gemini 0.256 0.408 0.644 0.052 0.104 0.206 0.203 0.014 0.024 0.049 0.076 0.030 0.128 0.221 0.308
Claude 0.201 0.422 0.658 0.014 0.105 0.269 0.281 0.011 0.033 0.079 0.128 0.087 0.069 0.183 0.277
DeepSeek-V3 0.190 0.401 0.621 0.001 0.090 0.260 0.217 0.001 0.027 0.078 0.105 0.013 0.102 0.246 0.314
DeepSeek-R1 0.224 0.447 0.651 0.001 0.197 0.463 0.496 0.005 0.024 0.068 0.096 0.024 0.148 0.326 0.414
Llama-3.1-70B 0.238 0.342 0.609 0.003 0.097 0.164 0.210 0.012 0.037 0.050 0.116 0.109 0.124 0.185 0.312
Book GPT-4o-mini 0.059 0.159 0.475 0.003 0.035 0.081 0.446 0.003 0.013 0.038 0.581 0.044 0.036 0.093 0.501
GPT-4o 0.088 0.192 0.567 0.027 0.057 0.133 0.472 0.021 0.011 0.024 0.500 0.445 0.052 0.116 0.513
Gemini 0.076 0.221 0.623 0.011 0.035 0.135 0.319 0.013 0.014 0.044 0.274 0.072 0.042 0.133 0.405
Claude 0.054 0.193 0.608 0.010 0.043 0.161 0.515 0.010 0.020 0.068 0.444 0.056 0.039 0.141 0.522
DeepSeek-V3 0.040 0.124 0.667 0.008 0.056 0.190 0.385 0.005 0.014 0.047 0.230 0.066 0.037 0.120 0.351
DeepSeek-R1 0.072 0.230 0.471 0.018 0.060 0.194 0.167 0.031 0.015 0.051 0.333 0.097 0.049 0.159 0.324
Llama-3.1-70B 0.073 0.170 0.542 0.022 0.038 0.082 0.470 0.052 0.014 0.028 0.178 0.158 0.042 0.093 0.397
  • Precision: The proportion of recommended items that are actually relevant to the user’s query.
  • Recall (He et al., 2020): The proportion of relevant items that are successfully recommended out of all relevant items available.
  • Condition Match Rate (CMR): The percentage of recommended items that meet the specified conditions. For Condition-based Queries, it is important to check if the model’s recommendations strictly align with the given conditions, as items that don’t meet these are likely unsatisfactory. Traditional metrics like Precision and Recall are insufficient for evaluating this alignment. Therefore, we propose CMR to achieve this goal.
  • Fail to Recommend (FTR): This metric measures the proportion of queries where the model fails to generate recommended items. A lower FTR is preferred, indicating the model can generate a list of items. However, in Misinformed Condition-based Query, a higher FTR shows the model’s ability to identify and reject incorrect information, thereby avoiding wrong recommendations.
3.2. Results on Condition-based Query
Observation 1. For different LLMs, GPT-4o and DeepSeek-R1 outperform other models. As shown in Table 2, GPT-4o achieves the highest Precision and the second-highest CMR average score on both the movie and book datasets, while DeepSeek-R1 ranks first in Recall and second in Precision. Notably, models with advanced reasoning capabilities, such as DeepSeek-R1, excel in handling query types that require reasoning. For example, in Implicit Condition Query, DeepSeek-R1 exhibits a significantly smaller drop in performance compared to other models when transitioning from explicit to implicit conditions. This positions it as the leader in Implicit Condition Query performance. It suggests that the advanced reasoning ability enables the LLM-based assistant to infer implicit attributes not explicitly mentioned in the conditions, leading to more accurate recommendations. There is a related case study in Appendix D.2.
Observation 2. For different query types, model performance decreases with increasing query difficulty. In both the movie and book datasets, most LLMs perform best on Explicit Condition Query, outperforming the other two query types. This suggests that LLMs are better at handling queries that clearly state conditions. In contrast, implicit conditions present a greater challenge because models must infer constraints about items that are not explicitly stated. It is worth noting that models with strong reasoning capabilities like Deepseek-R1 maintain strong performance, suggesting benefits from its reasoning capability. In addition, the misinformed condition proves most difficult, with the lowest recall and CMR, indicating frequent failures in providing personalized recommendations when queries contain misleading information. Moreover, higher FTR rates under this condition suggest that LLM-based assistants may leverage general knowledge to detect misinformation in queries and avoid directly generating inaccurate recommendations.
Figure 4. Performance on Condition-based Query with different number of conditions.
Observation 3. For different numbers of conditions in queries, Precision and Recall improve with more conditions, while CMR declines for Explicit Condition Query and rises for others. As Figure 4 illustrated, we take the movie dataset as an example and investigate the impact of varying the number of conditions in queries on model performance. We note that additional conditions narrow the recommendation scope, thus improving the recommendation accuracy. For example, the set of candidates for “Action movies by Nolan released after 2010” is much smaller than “movies by Nolan”, reducing ambiguity and improving the relevance of the recommendations. The impact of condition quantity on CMR metrics varies across query types. In Explicit Condition Query, CMR is highest with a single condition, as adding more conditions increases the difficulty of satisfying all constraints, leading to reduced coverage. In contrast, for Implicit and Misinformed Condition Query, additional conditions provide more context, allowing the model to better infer implicit conditions and reduce the impact of factual error in the conditions. As a result, the CMR metric improves with more conditions in these query types.
Observation 4. For the effect of user-item interaction history, incorporating them enhances recommendation quality in two key ways. The effect of incorporating user-item interaction history is shown in Figure 5. First, incorporating user-item history improves Precision across all query types by generating a more personalized recommendation list. The reason can be that queries with limited conditions (e.g., “movies directed by Nolan”) often yield a large candidate pool. By leveraging historical interactions, the model can prioritize select items that align with the user’s preferences and satisfy the query conditions, effectively filtering irrelevant candidates. We provide examples of both using and not using history in the Appendix D.3.
Figure 5. The effect of incorporating user-item history.
Second, incorporating user history does not always improve CMR. By analyzing the output of the LLM assistant, we note that adding history can also introduce “distractor” items that reduce the model’s strict condition adherence. For example, the LLM assistant may recommend “The Prestige” for “Nolan films” due to the user’s fantasy genre bias in history. This suggests that incorporating history may divert the model’s attention, making it more challenging to strictly adhere to the given conditions. These results highlight an important trade-off: while incorporating history can enhance the model’s understanding of user preferences, it may also hinder the model’s ability to closely align the specific requirements in the query. Consequently, balancing the influence of historical data with the need for precise condition adherence is crucial for optimizing performance in recommendation assistants.
3.3. Results on User Profile-based Query
Observation 5. For different LLMs, Gemini-1.5 Pro and DeepSeek-R1 demonstrated better performance compared to other models. As shown in Table 3, larger and more advanced models, such as DeepSeek-R1 and Gemini-1.5-Pro, consistently achieve higher Precision and Recall. In contrast, smaller models like GPT-4o-mini tend to lag behind. From the FTR (Failure to Recommend) metrics perspective, most models exhibit a low failure rate, indicating their ability to effectively follow instructions and generate recommendation lists in the required format.
Table 3. Performance on User Profile-based Query. The  /  means the best/the second-best result, respectively.
Domain Model Interest-based Query Demographics-based Query Average
[…]
Movie GPT-4o-mini 0.013 0.058 0.000 0.018 0.054 0.000 0.015 0.056 0.000
GPT-4o 0.018 0.067 0.001 0.021 0.059 0.000 0.020 0.063 0.000
Gemini 0.019 0.072 0.007 0.019 0.063 0.000 0.019 0.072 0.004
Claude 0.015 0.082 0.000 0.018 0.054 0.000 0.017 0.068 0.000
DeepSeek-V3 0.015 0.071 0.000 0.019 0.060 0.000 0.017 0.066 0.000
DeepSeek-R1 0.014 0.081 0.000 0.015 0.068 0.000 0.015 0.075 0.000
Llama-3.1-70B 0.014 0.061 0.000 0.015 0.046 0.000 0.015 0.054 0.000
Book GPT-4o-mini 0.038 0.104 0.004 - - - 0.038 0.104 0.004
GPT-4o 0.043 0.101 0.022 - - - 0.043 0.101 0.022
Gemini 0.056 0.127 0.049 - - - 0.056 0.127 0.049
Claude 0.018 0.072 0.012 - - - 0.018 0.072 0.012
DeepSeek-V3 0.020 0.081 0.005 - - - 0.020 0.081 0.005
DeepSeek-R1 0.030 0.112 0.031 - - - 0.030 0.112 0.031
Llama-3.1-70B 0.049 0.098 0.003 - - - 0.049 0.098 0.003
Observation 6. For different query types, Demographics-based Query exhibit lower Recall than Interest-based Query but similar Precision. Across all models, Recall for Demographics-based Query is generally lower than for Interest-based Query. This suggests that LLMs struggle to identify relevant recommendations when relying on user demographics. The difference in Recall can be attributed to the distinct ways in which these queries are constructed. Interest-based Query leverages clear interest, making it easier for models to generate relevant recommendations. In contrast, Demographics-based Query requires recommendation assistants to infer preferences from broader demographic and behavioral trends, which introduces more variability and ambiguity.
Figure 6. Impact of Interest Popularity on Precision & Recall
Observation 7. For Interest-based Query, queries constructed from less prevalent interests tend to be more challenging. As shown in Figure 6, we categorized the queries into five groups based on their popularity and conducted experiments on the movie and book datasets. It can be observed that interests that are shared by more users tend to yield higher Precision and Recall. This is likely because widely shared interests are more easily recognizable and interpretable by LLMs, leading to better performance. In contrast, the book domain shows an opposite trend. This is due to the nature of book data and the evaluation process. Books often have multiple editions or publishers, and popular titles tend to have more variants. Therefore, the popular book (e.g., “Dune”) can suffer lower metrics because LLMs confuse editions/publishers (e.g., “Dune: 1965” vs. “2021 adaptation”). For less popular books, fewer variants exist, making it easier for LLMs to provide exact matches, which leads to higher performance metrics.
Observation 8. For Demographics-based Query, the performances of LLMs show notable variations based on different user demographics. For instance, as figure 9 in the Appendix E shows, we note that models exhibit higher accuracy for female users. It reflects more consistent preference patterns of female users (e.g., stable genre preferences like romantic comedies), which allows the model to make accurate recommendations. Regarding occupation, the model performs best for users in sales/marketing roles, likely because these professions are associated with more consistent and recognizable patterns in user behavior. Age-wise, LLMs perform better regarding Recall for people aged 50–55. This is likely because their preferences are more focused and less affected by rapidly changing popular culture compared to younger users, who tend to have more diverse preferences. For users aged 56 and above, the model’s effectiveness is not as strong as it is for the 50–55 age group, possibly because older users engage less frequently in online activities, such as rating or reviewing movies. As a result, there is less data about this age group in the training set of LLMs, which makes it more challenging for models to understand their preferences.
4. Related Work
Recent advances in LLM-based recommender systems have led to the development of various benchmarks to evaluate their capabilities (Liu et al., 2023; Jiang et al., 2024; Tan et al., 2025; Xue et al., 2025; Zhang et al., 2023a). Early efforts like LLMRec (Liu et al., 2023) focus on traditional recommendation tasks (e.g., rating prediction, sequential recommendation), establishing baselines for LLMs against classical methods. In addition to traditional recommendation tasks, recent studies have increasingly focused on evaluating the unique characteristics of LLM for recommendations. For instance, one recent study (Jiang et al., 2024) introduces a multi-dimensional framework that assesses not only utility and novelty but also LLM’s specific behaviors like history length sensitivity (how recommendations perform with sparse user histories) and hallucination (generation of non-existent items). Similarly, PerRecBench (Tan et al., 2025) isolates personalization accuracy by eliminating biases in users and item ratings, testing whether LLMs infer user preferences.
Meanwhile, with the rise of LLMs, varying benchmarks have been proposed to evaluate their capabilities as assistants. For example, Gaia (Mialon et al., 2023) measures LLM’s capabilities as a general-purpose AI assistant. In addition, domain-specific frameworks like VoiceBench (for voice assistant) (Deng et al., 2024) and Mobile-Bench (for mobile assistant) (Chen et al., 2024) assess the ability of LLM as the specific assistant. However, despite progress in both recommendation and assistant evaluation, there remains a gap in assessing LLMs as personalized recommendation assistants. To bridge this gap, we introduce RecBench+, a novel benchmark designed to evaluate LLMs’ potential as personalized recommendation assistants. Due to the space limitation, some studies about RecSys studies are reviewed in Appendix A.
5. Conclusion
In this paper, we introduce a new paradigm for next-generation recommender systems and propose a novel benchmark dataset, RecBench+, to evaluate the potential of LLMs as personalized recommendation assistants. Our benchmark includes five specially designed query types that comprehensively assess LLM-based assistants’ capabilities in handling complex user queries in real-world scenarios. We conduct extensive experiments on a diverse set of LLMs, including those with advanced reasoning capabilities. Through these experiments, we present eight detailed observations that shed light on the strengths and limitations of LLMs in serving as personalized recommendation assistants. We hope our benchmark and findings lay a solid foundation for future research and development in this emerging domain. We provide more discussion about future directions in Appendix B.
Appendix A More Related Work
RecSys as intelligent filtering tools have gained rapid development in the past decades. Early methods such as NCF (He et al., 2017) and SVD++ (Koren, 2008) leverage collaborative filtering (CF) information to generate recommendations. With the evolution of deep neural networks, Graph Neural Networks (GNNs) have dominated RecSys due to their ability to model high-order users’ CF information (He et al., 2020; Fan et al., 2019; Ying et al., 2018). Typical studies like LightGCN (He et al., 2020) and GraphRec (Fan et al., 2019) introduce GNN to aggregate neighborhood information for user and item embeddings to improve recommendations.
Recently, with the emergence of LLMs, various LLM-based RecSys have attracted widespread attention. With powerful reasoning and language understanding capabilities, LLMs can delve deeper into the semantic understanding of user preferences and item characteristics (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023b; Qu et al., 2024; Wang et al., 2025). For example, P5 (Geng et al., 2022) unified various recommendation tasks as the text generation task. TALLRec (Bao et al., 2023) and CoLLM (Zhang et al., 2023b) further tuned on user-item interaction data to compensate for inadequate recommendation data during the pre-training phase. Although existing recommendation frameworks can address conventional tasks like rating prediction (Geng et al., 2022; Koren, 2008) and sequential recommendation (Bao et al., 2023; Zhang et al., 2023b; Wang et al., 2025), their ability to handle diverse, complex, and multi-faceted user needs remains an open question. Most of them directly transform the datasets from traditional recommendation paradigm to the interactive paradigm in the era of LLMs, which is inappropriate for practical scenarios and insufficient for evaluating recommendation performance. Note that modern users increasingly require personalized and context-aware recommendations that deeply understand multifaceted user needs. To fill this research gap, we propose a holistic benchmark to systematically evaluate the potential of various advanced LLMs to serve as recommendation assistants.
Appendix B Future Directions
In this study, we proposed a novel paradigm for evaluating the LLM-based recommendation. We systematically examined the potential of widely used LLM backbones to serve as personalized recommendation assistants in our RecBench+ benchmark dataset. Our findings showed that LLMs already exhibit preliminary capabilities as personalized recommendation assistants, yet their full potential remains under-explored. This highlights the need for further investigation to fully unlock the power of LLMs in personalized recommendation tasks. Several future directions present exciting opportunities to enhance and expand the capabilities of LLM-based recommendation assistants.
One important direction to explore is the effectiveness of LLMs in real-world recommendation scenarios, especially after specialized fine-tuning. Our study shows that pure LLM backbones possess initial capabilities as personalized recommendation assistants, and different LLMs exhibit varying strengths in handling specific types of user queries. For instance, models like GPT-4o and DeepSeek-R1 demonstrate strong performance in handling user queries with explicit conditions for expected items, whereas Gemini-1.5-Pro excels in processing queries without explicit conditions, requiring user profile understanding. Therefore, it is necessary to explore how specialized fine-tuning can further optimize their performance as personalized recommendation assistants. This includes evaluating whether such training can lead to improvements across different tasks, the extent of these improvements, and whether certain models benefit more than others from specialized training.
Another key direction for improvement is integrating external tools to enhance the performance of LLM-based recommendation assistants. External tools, such as search engines and knowledge bases, can provide valuable up-to-date information and real-world context that LLMs may lack due to the static nature of their training data. By incorporating these tools, recommendation assistants can better understand the world and current events, enabling them to provide more timely, intelligent, and customized recommendations. This integration could greatly enhance the alignment of recommendations with users’ evolving needs and preferences, delivering more personalized and intelligent recommendations.
Finally, in domains where LLMs lack specialized knowledge—such as in e-commerce recommendations—combining assistants with techniques like Retrieval-Augmented Generation (RAG) or other hybrid models could prove highly effective. LLMs typically do not have access to domain-specific knowledge, such as detailed item pools or up-to-date product information, making them ill-equipped to provide accurate recommendations in these areas. RAG techniques may address this limitation by using external retrieval mechanisms to supplement the generative capabilities of LLMs. This allows the model to dynamically query databases or corpora of item information, enhancing the accuracy and contextual relevance of the recommendations. Without such integration, LLMs alone may struggle to offer satisfactory recommendations in specialized domains.
Appendix C Query Example
Table 4. Examples of queries from different types in our RecBench+ benchmark.
Task Query
Explicit Condition Query I’m really interested in classic films and would love to watch something that showcases Charlie Chaplin’s legendary comedic talent. Additionally, I’ve heard that Roland Totheroh’s cinematography adds an exceptional visual quality to movies. If you could point me in the direction of films that include both of these elements, I’d greatly appreciate it!
Implicit Condition Query I recently watched Clockers (1995) and Bamboozled (2000), and I was really impressed by the direction in both films. I’m eager to explore more works from the director, as I found their storytelling style and vision very engaging. If you could suggest other films associated with this director, that would be fantastic.
Misinformed Condition Query I recently watched Lorenzo’s Oil and was really impressed by the cinematography done by Mac Ahlberg. I’m interested in finding more films that showcase his cinematographic style. I also remember seeing his work in Beyond Rangoon, so if there are any other movies he contributed to, I’d love to check them out!
Interest-based Query I’m fond of romantic and dramatic films from the golden age of Hollywood like ‘Roman Holiday’ and ‘My Fair Lady’. Are there any other dramatic romances from that period you would recommend?
Demopraphics-based Query I’m a psychology professor and I’m looking for movies that delve into human emotions and relationships. Have you got any?
Examples of different categories of queries in our benchmark are shown in Table 4.
Appendix D Case Study
We carefully examined LLM-based recommendation assistants’ outputs and identified three key factors significantly influencing the recommendation performance. In the following sections, we provide examples of each of these factors.
D.1. The Impact of LLM Knowledge on Performance
Model knowledge plays a crucial role in the quality and relevance of recommendations. In our evaluation, we observed that recommendation assistants often fail to generate suitable recommendations when they lack sufficient domain-specific knowledge, even if the query explicitly states conditions for the user’s expected items.
In the case presented, GPT-4o-mini fails to recommend movies featuring cinematography by Ellery Ryan, responding that it lacks the necessary information. In contrast, GPT-4o successfully recommends relevant films, such as I Love You Too (2010) and The Rage in Placid Lake (2003). This performance difference suggests that GPT-4o may possess a more extensive knowledge base or improved knowledge-recalling mechanisms, allowing it to generate appropriate recommendations where the other fails.
D.1 Case for Impact of LLM Knowledge Query: I’m looking for a movie with stunning visuals, specifically those that have the unique touch of cinematography by Ellery Ryan. I’ve heard that his work enhances storytelling, and I’d love to experience it firsthand. Output (GPT-4o-mini): I’m sorry, but I am unable to provide movie recommendations specifically featuring cinematography by Ellery Ryan, as I do not have data on his work. Output (GPT-4o): I Love You Too (2010), The Rage in Placid Lake (2003)
D.2. The Impact of Advanced Reasoning Capability on Performance
Advanced reasoning is critical in improving the robustness and accuracy of recommendation assistants, especially when a user’s query contains factual errors that could serve as a distraction. LLMs with strong reasoning capabilities are more likely to verify information, resolve inconsistencies, and generate appropriate recommendations.
In the first case study, the user inquires about movies featuring cinematography by John Blick, referencing The Mirror Has Two Faces (1996), which is incorrect. Recommendation assistants lacking a clear reasoning process would attempt to match John Blick with known cinematographers, leading to incorrect recommendations. However, DeepSeek-R1, which possesses advanced reasoning capability, successfully identifies that John Blick was not the cinematographer of the mentioned film, instead recognizing the correct individual, Dante Spinotti. By reasoning through the discrepancy, the model does not give any movies directed by John Blick.
D.2 Case 1 for Impact of Advanced Reasoning Capability Query: I recently watched The Mirror Has Two Faces and was really impressed by John Blick’s cinematography. I’m interested in finding more films that feature his work behind the camera. If you know of any other movies that he did the cinematography for, I’d love to check them out! Ground Truth: Falling Down (1993), Speed (1994), The Devil’s Advocate (1997), U.S. Marshals (1998); or ”None” Output (DeepSeek-V3): The Last Seduction, One False Move, The Underneath, Albino Alligator, The Lookout Output (DeepSeek-R1): None Reasoning (DeepSeek-R1): … first, I need to confirm if John Blick is the actual cinematographer for that movie. Let me check my database. Ah, T̈he Mirror Has Two Faces(̈1996) was actually cinematographed by Dante Spinotti, not John Blick. So there’s a mistake here. Maybe the user misremembered the name. Alternatively, maybe John Blick is a lesser-known DP or a pseudonym? Let me cross-reference. No, seems no record of John Blick working on that film…Well, given the instructions, I need to answer with the movies, no explanations. So since there are none, the field is empty.
D.2 Case 2 for Impact of Advanced Reasoning Capability Query: I recently enjoyed the cinematography in Taps (1981) and Absence of Malice (1981), both of which showcased Scott Ambrozy’s incredible work. I’m eager to find more films that feature his cinematography style. Any recommendations? Ground Truth: 3 Days of the Condor (a.k.a. Three Days of the Condor) (1975), Network (1976), The Exorcist (1973), The Stepford Wives (1975) ; or ”None” Output (DeepSeek-V3): The Chosen (1981), Four Friends (1981), Split Image (1982), Heart Like a Wheel (1983) Output (DeepSeek-R1): The French Connection (1971), The Exorcist (1973), Network (1976), Three Days of the Condor (1975), The Electric Horseman (1979), Straight Time (1978), The Prince of the City (1981), Wyatt Earp (1994). Note: The films listed above feature the cinematography of Owen Roizman, who shot Taps and Absence of Malice. Scott Ambrozy appears to be a confusion or typo in the query. Reasoning (DeepSeek-R1): … Wait, first, who is Scott Ambrozy? I know some big cinematographers, but this name isn’t ringing a bell. Maybe I should verify if there’s a Scott Ambrozy in the industry. Let me think. Maybe the user misspelled the name? Looking up Taps on IMDb, yeah, Owen Roizman is the cinematographer for Taps. Absence of Malice also lists Owen Roizman. So Scott Ambrozy might be a confusion with another name. So the user probably meant Owen Roizman. So to answer correctly, I should recommend other films by Owen Roizman. Since the user wants movies with his cinematography style, I should list those. Roizman’s notable works include…
A similar pattern emerges in the second case, where the user asks for films by Scott Ambrozy, citing Taps (1981) and Absence of Malice (1981). The user is actually seeking recommendations for movies with the same cinematographer as these two films. However, the model DeepSeek-V3, lacking strong reasoning capabilities, tries to list movies by Scott Ambrozy, resulting in incorrect recommendations. In contrast, DeepSeek-R1 takes a different approach. It questions whether Scott Ambrozy is actually a cinematographer first. Subsequently, it identifies that the actual cinematographer of these films is Owen Roizman. Finally, DeepSeek-R1 infers that the user probably meant to ask about Roizman’s work, providing correct recommendations.
These examples highlight that the ability to evaluate and refine its reasoning process significantly enhances recommendation accuracy. Instead of passively following instructions, models with advanced reasoning capabilities proactively identify inconsistencies, validate claims, and generate factually sound responses. This improvement is particularly impactful in real-world applications, where user inputs often contain typos, misinformation, or ambiguous phrasing.
D.3. The Impact of User-Item Interaction History on Performance
User-item interaction history is crucial for personalized recommendation assistants to align their recommendation outputs with individual user preferences. Assistants lacking access to historical data may generate recommendations that meet the user’s criteria yet fail to fully align with their preferences. Our case study demonstrates how historical preferences influence recommendation accuracy.
D.3 Case for Impact of User-Item Interaction History Query: I’m looking for a nostalgic film experience, ideally featuring the talented duo of Karen Dotrice and Matthew Garber. Their performances always bring a sense of charm and whimsy that I adore, so I’d appreciate any recommendations that showcase them together! User History: Starship Troopers (1997), You’ve Got Mail (1998), Beetlejuice (1988), Pocahontas (1995), Beauty and the Beast (1991), Rescuers, The (1977), On Golden Pond (1981), Star Wars: Episode V - The Empire Strikes Back (1980), Bowfinger (1999) Ground Truth: Mary Poppins (1964), The Gnome-Mobile (1967) Output (GPT-4o, w/o history): The Three Lives of Thomasina Output (GPT-4o, with history): Mary Poppins (1964), The Gnome-Mobile (1967) Note: *Mary Poppins* (1964) is a strong recommendation as it blends Disney’s charm with whimsical fantasy, aligning with the user’s preference for animated classics like *Beauty and the Beast* and *The Rescuers*. Its mix of adventure and magic also echoes the fantastical elements of *Beetlejuice* and the nostalgia of *You’ve Got Mail*. *The Gnome-Mobile* (1967) is another fitting choice, offering a lighthearted adventure with Karen Dotrice and Matthew Garber, while its humor parallels *Bowfinger*. In contrast, *The Three Lives of Thomasina* (1963) leans more toward sentimentality and drama, deviating from user’s favored whimsical and adventurous films.
When the LLM (GPT-4o) doesn’t have access to the user’s history, it recommends The Three Lives of Thomasina (1963). This movie features Karen Dotrice, but it’s more about sentimentality and drama. The recommendation misses the user’s preference for whimsical, adventurous, and nostalgic films. However, when user history is incorporated, the LLM-based assistant correctly recommends Mary Poppins (1964) and v. These movies match the user’s requirements for Karen Dotrice and Matthew Garber’s performances and fit the themes and styles of the films the user has enjoyed before.
This performance gap underscores the importance of incorporating past interactions into recommendation assistants. Without access to history, assistants mainly rely on the conditions for items in the query to provide recommendations, which can result in recommendations that meet the basic criteria but don’t fully satisfy the user’s preferences. In contrast, integrating user history allows personalized recommendation assistants to infer underlying preferences and provide recommendations that are not only conditionally relevant but also more aligned with user preferences. By analyzing past interactions, the assistant can refine its understanding of user preferences, leading to more personalized and contextually relevant suggestions. This highlights the importance of leveraging historical data to enhance recommendation accuracy and user satisfaction.
Table 5. Results on Condition-based Query with K=5
Model Explicit Condition Implicit Condition Misinformed Condition Average
Precision Recall CMR FTR Precision Recall CMR FTR Precision Recall CMR FTR Precision Recall CMR
GPT-4o 0.233 0.436 0.657 0.002 0.123 0.245 0.266 0.003 0.024 0.052 0.077 0.100 0.127 0.244 0.333
Gemini 0.223 0.416 0.607 0.002 0.099 0.205 0.195 0.003 0.023 0.048 0.062 0.003 0.115 0.223 0.288
Claude 0.234 0.436 0.647 0.002 0.126 0.247 0.277 0.010 0.035 0.075 0.105 0.063 0.132 0.253 0.343
Figure 7. Performances on Condition-based queries with different number of conditions when K=5.
Figure 8. The effect of incorporating user-item history when K=5.
Appendix E Visualization of Performance of Queries Constructed based on Different User Demographics
For Demographics-based Queries, the performances of LLMs show notable variations based on different user demographics. We visualize the trends in figure 9.
Figure 9. Average Recall of queries constructed based on different user demographics.
Appendix F Experiments with Fixed K for Recommendation Performance
We evaluated the performance of several well-performing models by testing their ability to generate top K recommended items. The prompt used for this experiment involved modifying the one in  G from “Please output the recommended movies directly” to “Please output {K} recommended movies directly.” Table  5 presents the experimental results on the movie dataset. The findings are consistent with the observations in the main text: the LLMs perform best on Explicit Condition Query, and their performance gradually decreases on Implicit Condition and Misinformed Condition Query.
Furthermore, we analyzed the relationship between model performance, the number of conditions, and user history. The results are illustrated in Figures 7 and 8, and the observed trends align with the observations in the main text.
Appendix G Prompts for Model Testing
The prompts we use for testing different LLMs are:
  • w/o user-item interaction history: You are a movie recommender system. You are given a query from the user. You need to recommend movies that satisfy the user’s query. Please output the recommended movies directly; do not output other content.
  • with user-item interaction history: You are a movie (/book) recommender system. You are given a user’s watched movies (/reading history) and a query from the user. You need to recommend movies (/books) that satisfy the user’s query and user preferences. Please output the recommended movies directly; do not output other content.
Appendix H Prompts for Query Construction
The prompts we use for constructing queries with LLMs are shown in Table LABEL:tab:query_generation_prompt.
Table 6. Summarization of prompts for query generation
Task Prompt
Explicit Condition Query You are given a set of attributes. Please simulate a real user and generate a natural language query covering these attributes to search or request recommendations for related movies. The attributes are {Explicit Conditions} …
Implicit Condition Query You are given a set of attributes and relevant information for [MASK] attributes. Your task is to generate a query that meets the following criteria: - Query should ask for items that share the input attributes. - Do not directly mention the [MASK] attribute that has additional relevant information (e.g., name of cinematography or starring role) in queries. Instead, describe the [MASK] attribute using the relevant information provided. The input attributes and relevant information is: …
Misinformed Condition Query You are given a set of attributes and relevant information for one of the attributes. Your task is to generate query that meets the following criteria: - Imagine you are a real user, queries should ask for movies that share the input attributes. - One of the attributes is provided with relevant information, such as the movie name, the relation between the movie and the attribute, and the person’s name. You should describe this attribute using the relevant information provided. - There may be factual errors between the relevant information and the corresponding attribute, but you still need to describe it based on the relevant information provided without making corrections or explanations. The input attributes and relevant information is: …
Interest-based Query ## Input Introduction You are given a “Popular Movie List” and “Previous Movie Statistics”. “Popular Movie List” represents movies collectively watched by multiple users, while ”Previous Movie Statistics” refers to the statistics of movies that some users watched before watching the “Popular Movie List”. Here is an example: - Popular Movie List: [A, D, M, …] - Previous Movie Statistics: - [Z, V, K], count: 6 - [M, O, Z,Y], count: 5 - [Y, C, E, O, Z], count: In this example, 6 users watched [A, D, M, …] after watching [Z, V, K] ## Task Introduction 1. **Step 1: Generate Reasons** Please use all your knowledge about these movies to analyze why users watch movies in the Popular Movie List after watching Previous Movies. What is the relationship between them? Do not give reasons that are too far-fetched or broad. If you think it cannot be explained, don’t force an explanation. 2. **Step 2: Generate User Queries and Answers** For each reason you generate: - Create **realistic user queries**. These queries should simulate how a real user might ask recommender systems for movie recommendations using natural language. The query should be as complex and rich as possible, close to the reason, and specific enough not to make the recommendation system recommend movies other than the subset of answers. - Provide a **subset of the Popular Movie List/Previous Movies** as the answer, ensuring the recommendations align with the reason and query. This subset will serve as the answer to the query, so do not mention the movie title of the subset in the query. ## Output Format The output must follow this JSON structured format: [ ‘reason ’: , ‘query ’: , ‘movie subset ’: , ‘reason ’: , ‘query ’: , ‘movie subset ’: , … ]
Demographics-based Query You are tasked with analyzing a specific user group and their movie preferences. Based on the provided data, you should generate structured reasons, user queries, and answers in a clear and organized format. Follow these steps to complete the task: ## **Input Data**: 1. **User Group**: A description of the user group. 2. **List of Movies**: A ranked list of movies relevant to the user group, including their titles, release years, TF-IDF scores, and the percentage of users in the group who viewed them. ## **Your Task**: 1. **Step 1: Generate Reasons** Analyze the user group description and the provided movie list. Generate a structured list of reasons explaining why this user group prefers these movies. 2. **Step 2: Generate User Queries and Answers** For each reason you generate: - Create **realistic user queries** that naturally combine the reason and user group characteristics. These queries should simulate how a real user from the group might ask recommender systems for movie recommendations using natural language. - Provide a **subset of the movie list** as the answer, ensuring the recommendations align with the reason and query. Use the movie title to represent a movie.